OpenStudy (anonymous):

A light on a computer comes on for 26,700 microseconds. One microsecond is 10^-6 seconds. Work out the length of time, in seconds, that the light is on (a) in standard form and (b) as a decimal.

OpenStudy (anonymous):

1 microsecond = 10^-6 seconds 26700 microseconds = 26700* 10^-6 seconds

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

Yes, thanks again! :)

OpenStudy (anonymous):

Did u convert it into standard form and decimal?

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

GREAT

OpenStudy (anonymous):

If one microsecond is equal to \[10^{-6}\] then, 26700 microseconds are equal to: \[26700 \times 10^{-6} \] and 26700 can also be written as: \[2.67 \times 10^{4}\] so your whole equation would be: \[2.67 \times 10^{4} \times 10^{-6}\] which according to the exponential rules is: \[2.67 \times 10^{4-6}\] which is: \[2.67 \times 10^{-2}\] which in decimal form is:\[0.0267\]

OpenStudy (anonymous):

Thanks loads! :)